# FineFake: A Knowledge-Enriched Dataset for Fine-Grained Multi-Domain Fake News Detection

### v2

@@ 6.1. Baseline Models
Following previous works (zhou2020survey, ), we select the following methods as baselines and divide them into three categories: content-based method, knowledge-enhanced method and multi-domain method.
@@
Content-Based Multimodal Method. MVAE (khattar2019mvae, ) comprises three components: an encoder to encode the shared representation of features, a decoder to reconstruct the representation, and a detector to classify the truth of posts. SAFE (Zhou2020SAFESM, ) calculates the relevance between textual and visual information and defines it as cosine similarity modification to detect fake news.
Knowledge-Enhanced Method. CompNet (hu2021compare, ) constructs a directed heterogeneous document graph to utilize knowledge base. KAN (dun2021kan, ) incorporates semantic-level and knowledge-level representations in news to improve the performance for fake news detection. KDCN (sun2023inconsistent, ) captures two level inconsistent semantics in one unified framework to detect fake news.
Multi-Domain Method. EANN (wang2018eann, ) firstly utilizes a discriminator to derive event-invariant features for multi-domain fake news detection. MDFEND (nan2021mdfend, ) introduces a multi-domain fake news detection model that leverages a domain gate to aggregate multiple representations extracted by a mixture of experts. M3FEND (zhu2022memory, ) proposes a memory-guided multi-view framework to address the problem of domain shift and domain labeling incompleteness. CANMD (yue2022contrastive, ) proposes a contrastive adaptation network to solve the label shift problem in early misinformation detection.
6.2. Implementation Details
@@
All experiments are conducted using the proposed FineFake dataset. In classification experiments, instances within FineFake are split into the ratio of 6:2:2 for training, evaluating and testing, respectively. In domain adaptation experiments, source domain data are split into the ratio of 9:1 for training and evaluating, while all data from target domain are used for testing. The experiments are executed on a computational setup consisting of 4 NVIDIA 3090 GPUs, each equipped with 24GB of memory. We set $\alpha$ as 0.8 and $\beta$ as 0.2 for loss function. For optimization purposes, we employ the AdamW optimizer (loshchilov2017decoupled, ) with a weight decay value of 5e-4. The batch size is set to 32, and the initial learning rate is established at 1e-3. Subsequently, the learning rate was decayed gradually with each epoch. In order to eliminate the potential impact of random variations, we fix the random seed throughout the experiments. The hyperparameters of all baseline models are carefully tuned on the validation sets to achieve an optimal configuration.
@@
Table 6. Multi-Domain task results, best results are in bold and second best results are underlined. In Task 1, we experiment with entertainment and conflict topics as tests, using the remaining four topics for training on the same platform.
@@
Target Enter. Conflict Stream Official Fact-Check Stream Official Fact-Check
@@
6.3. Experimental Results
6.3.1. Binary Classification Task
In this study, we perform binary classification analysis under two conditions: knowledge enhancement and no knowledge. Experimental results are presented in Table 5. Notably, KDCN and KEAN exhibit the highest improvement when trained on the knowledge-augmented dataset, indicating their robust capacity to assimilate and leverage external knowledge. Across all knowledge-enhanced models, training with knowledge enhancement yields a substantial improvement in all four metrics. This finding demonstrates the significance of incorporating high-quality external knowledge, which further reveals the value of knowledge provided in the proposed FineFake dataset.
6.3.2. Fine-grained Classification Task
In this task, we utilize the fine-grained annotations for both training and testing purposes. We employ accuracy, macro recall, macro precision and macro f1-score to assess the performance of baseline models. Experimental results are presented in Table 5. Notably, there is significant degradation in the performance of all methods on fine-grained classification. This experiment enables more detailed classification of fake news, thereby enhancing the interpretability of fake news detection.
6.3.3. Multi-domain Adaptation Task
@@
Here we evaluate the performance of different models under the multi-domain adaption scenario by the three domain adaptation tasks defined in 3.4. Table 6 presents the experimental results. Given significant disparities in the percentage of positive and negative samples across certain domains, e.g. Official Stream, the weighted f1 value is adopted as the evaluation index for all experiments.
@@
In the topic domain adaptation task, we select entertainment and conflict topic as the test data, while utilizing data from the remaining four topics for training. Results demonstrate varying degrees of content differences between topics. Notably, all model’s experience a substantial decline in performance from the entertainment topic to the conflict topic, indicating that the conflict topic exhibits greater variance in data distribution. This finding aligns with our analysis of FineFake in 3.4. KEAN achieves the most SOTA metrics in this task, showcasing its adaptability across different topics. KEAN reaches most SOTA metrics in this task, which demonstrates our model exhibits good adaptability across different topics.
@@
In platform domain adaptation task, we choose politics data as the training and testing data. Politics is chosen due to its high proportion among the six topics and its significant attention in research (wang2017liar, ; roy2019deep, ; krstovski2022evons, ). KEAN achieves most SOTA in both metrics, particularly when source and target datasets exhibit significant label shift problems, e.g fact-check source domain and official news test domain. This is attributed to our utilization of re-weighting methods, enabling our model to possess superior generalization capabilities when encountering label shift issues.
@@
In the dual domain adaptation task, models are trained with conflict topic data from the source platform and tested with politics topic data from the target platform. Results show a significant decrease in performance across all models on this task, as crossing two domains exacerbates both covariate shift and label shift problems. The pronounced domain shift poses a formidable challenge to the model’s generalization capability. For instance, when encountering substantial label and convariate shift between two news environments(official conflict data with stream politics data), models lose ability to distinguish real from fake news. This indicates the benchmark role of FineFake in future research on fake news detection, offering enhanced generalization capability and practical value. However, KEAN still reaches SOTA in most task settings, further confirming its cross-domain capability to address convariate shift and label shift simultaneously.
Figure 4. Case studies on our dataset FineFake.
6.4. Case Study
@@
7. Conclusion and Future Work

### v3

@@ VI-A Baseline Models
Following previous works [1], we select the following fake news detection methods as baselines and divide them into three categories: content-based method, knowledge-enhanced method and multi-domain method.
Content-Based Multimodal Method.
@@
  • MVAE [9] comprises three components: an encoder to encode the shared representation of features, a decoder to reconstruct the representation, and a detector to classify the truth of posts.
  • SAFE [30] calculates the relevance between textual and visual information and defines it as cosine similarity modification to detect fake news.
Knowledge-Enhanced Method.
  • CompNet [16] constructs a directed heterogeneous document graph to utilize knowledge base.
  • KAN [14] incorporates semantic-level and knowledge-level representations in news to improve the performance for fake news detection.
  • KDCN [33] captures two level inconsistent semantics in one unified framework to detect fake news.
Multi-Domain Method.
  • EANN [36] firstly utilizes a discriminator to derive event-invariant features for multi-domain fake news detection.
  • MDFEND [17] introduces a multi-domain fake news detection model that leverages a domain gate to aggregate multiple representations extracted by a mixture of experts.
  • M3FEND [35] proposes a memory-guided multi-view framework to address the problem of domain shift and domain labeling incompleteness.
  • CANMD [37] proposes a contrastive adaptation network to solve the label shift problem in early misinformation detection.
VI-B Implementation Details
@@
All experiments are conducted using the proposed FineFake dataset. In classification experiments, instances within FineFake are split into the ratio of 6:2:2 for training, evaluating and testing, respectively. In domain adaptation experiments, source domain data are split into the ratio of 9:1 for training and evaluating, while all data from target domain are used for testing. The experiments are executed on a computational setup consisting of 4 NVIDIA 3090 GPUs, each equipped with 24GB of memory. We set $\alpha$ as 0.8 and $\beta$ as 0.4 for loss function. For optimization purposes, we employ the AdamW optimizer [55] with a weight decay value of 5e-4. The batch size is set to 32, and the initial learning rate is established at 1e-3. Subsequently, the learning rate is decayed gradually with each epoch. In order to eliminate the potential impact of random variations, the random seed is fixed throughout the experiments. The hyperparameters of all baseline models are carefully tuned on the validation sets to achieve an optimal configuration.
VI-C Experimental Results
VI-C1 Binary Classification Task
In this study, we perform binary classification analysis under two conditions: knowledge enhancement and no knowledge. Experimental results are presented in Table V. Notably, KDCN and KEAN exhibit the highest improvement when trained on the knowledge-augmented dataset, indicating their robust capacity to assimilate and leverage external knowledge. Across all knowledge-enhanced models, training with knowledge enhancement yields a substantial improvement in all four metrics. This finding demonstrates the significance of incorporating high-quality external knowledge, further revealing the value of knowledge provided in FineFake dataset.
VI-C2 Fine-grained Classification Task
Accuracy, macro recall, macro precision and macro f1-score are employed in this task. Experimental results are presented in Table V. Notably, there is significant degradation in the performance of all methods on fine-grained classification. This indicates that the current models are not sufficiently effective in identifying the underlying causes of fake news, suggesting substantial room for improvement in fine-grained classification. KEAN’s superior performance may be due to the utilization of both images and knowledge, which enhances the model’s capability to comprehend the underlying reasons for falsehoods.
@@ VI-C2 Fine-grained Classification Task
TABLE VI: Multi-domain task results, best results are in bold and second best results are underlined. In Task 1, we experiment with entertainment and conflict topics as tests, using the remaining four topics for training on the same platform.
@@
Target Enter. Conflict Social Media Official Snopes Social Media Official Snopes
@@
VI-C3 Multi-domain Adaptation Task
@@
Here we evaluate the performance of models under the multi-domain adaption scenario by the three domain adaptation tasks defined in Section III-D. Table VI presents the experimental results. Given significant disparities in the percentage of positive and negative samples across certain platforms, e.g. official news with social media, the weighted f1 value is adopted as the evaluation index for all experiments.
@@
In topic domain adaptation task, the entertainment and conflict topic are selected as the test data, while utilizing data from the remaining four topics for model training. The results reveal notable variations in model performance across the different topics. Notably, all model’s performance is substantially declined from the entertainment topic to the conflict topic, indicating that the conflict topic exhibits greater variance in data distribution. This finding aligns with our analysis of FineFake in Section III-D. KEAN achieves most SOTA metrics in this task, showcasing its robustness and adaptability across different topics. Specifically, its ability to leverage external knowledge and cross-domain learning enables it to bridge the gap of distribution shift between training and testing data more effectively.
@@
In platform domain adaptation task, politics data is chosen due to its high proportion among the six topics and its significant attention in research [6, 56, 18]. Notably, there is a significant performance drop in all models during platform domain adaptation task, which can be attributed to the label shift between source and target platforms, leading to a higher likelihood of misclassification. KEAN achieves most SOTA in both metrics, particularly when source and target datasets exhibit significant label shift problems, e.g. Snopes to official news. This is attributed to our utilization of re-weighting methods, enabling our model to possess superior generalization capabilities when encountering label shift issues.
@@
In dual domain adaptation task, models are trained with conflict topic data from the source platform and tested with politics topic data from the target platform. Results show a significant decrease in performance across all models on this task, as crossing two domains exacerbates both covariate shift and label shift problems. The pronounced domain shift poses a formidable challenge to the model’s generalization capability. For instance, when encountering substantial label and convariate shift between two news environments (e.g. official conflict data with social media politics data), models lose ability to distinguish real from fake news. This indicates the benchmark role of FineFake in future research on fake news detection, offering enhanced generalization capability and practical value. KEAN still reaches SOTA in most task settings, further confirming its cross-domain capability to address convariate shift and label shift simultaneously.
Figure 5: Hyper-parameter sensitivity analysis of $\alpha$ and $\beta$.
VI-D Parameter Sensitive Analysis
We perform a hyper-parameter sensitivity analysis on two key parameters: $\alpha$ and $\beta$ in Formula 7, denoting the weights of the adversarial loss $\mathcal{L}_{adv}$ and the reconstruction loss $\mathcal{L}_{recon}$, respectively. This analysis is conducted in the context of the topic adaptation task on social media platforms, with the results presented in Fig 5. Notably, the model’s performance under conflict topic is more susceptible to the values of $\alpha$ and $\beta$, possibly due to greater divergence between the data in the conflict topic and the training data.
When both $\alpha$ and $\beta$ are set too low, the model struggles to learn sufficient domain-invariant knowledge, which is critical for generalization across different topics. As a consequence, its performance in cross-topic fake news detection declines significantly. On the other hand, when $\alpha$ and $\beta$ are set too high, the model becomes overly focused on learning domain-invariant representations, which diminishes its ability to perform the primary task of fake news detection. This excessive emphasis on domain adaptation leads to a reduction in task-specific learning, ultimately degrading the model’s overall performance. Balancing these two parameters is crucial to ensure the model captures both domain-invariant features and task-specific information effectively.
VI-E Case Study
Figure 6: Case studies on proposed dataset FineFake.
@@ VI-E Case Study
VII Conclusion and Future Work